# KKL Observer Synthesis for Nonlinear Systems via Physics-Informed Learning

### v1

@@ Abstract
This paper proposes a novel learning approach for designing Kazantzis-Kravaris/Luenberger (KKL) observers for autonomous nonlinear systems. The design of a KKL observer involves finding an injective map that transforms the system state into a higher-dimensional observer state, whose dynamics is linear and stable. The observer’s state is then mapped back to the original system coordinates via the inverse map to obtain the state estimate. However, finding this transformation and its inverse is quite challenging. We propose to sequentially approximate these maps by neural networks that are trained using physics-informed learning. We generate synthetic data for training by numerically solving the system and observer dynamics. Theoretical guarantees for the robustness of state estimation against approximation error and system uncertainties are provided. Additionally, a systematic method for optimizing observer performance through parameter selection is presented. The effectiveness of the proposed approach is demonstrated through numerical simulations on benchmark examples and its application to sensor fault detection and isolation in a network of Kuramoto oscillators using learned KKL observers.
@@
thanks: This work is supported by the Swedish Research Council and the Knut and Alice Wallenberg Foundation, Sweden. It has also received funding from the European Union’s Horizon Research and Innovation Programme under Marie Skłodowska-Curie grant agreement No. 101062523. Corresponding author: M. U. B. Niazi.
@@ 1 Introduction
Observers play a crucial role in applications such as output feedback control [54], fault diagnosis [16], and digital twins [17]. Accurate state estimation is critical for effective control, monitoring, and decision-making in various engineering domains. Full-state measurement of real-world systems is often impractical, or even impossible, due to limitations in sensing resources and capabilities. This necessitates the use of observers that employ available sensor measurements and the system’s model to estimate the state. A variety of observer designs exist, including Luenberger-like observers, extended Kalman filters, and high-gain observers, each with its own strengths and weaknesses depending on the specific application and system characteristics [24, 30, 7, 11]. This work focuses on Kazantzis-Kravaris/Luenberger (KKL) observers due to their applicability to a very general class of nonlinear systems and the well-developed theoretical framework for their design and analysis [29, 4, 2, 9, 12].
@@
Despite the advancements, synthesizing KKL observers remains a challenging task. The existing techniques assume the knowledge of an injective map that transforms a general nonlinear system into a special form required by the KKL observer. However, analytically finding this map is an inherently difficult problem. Moreover, in cases when this transformation map is known [8], finding its inverse so that one can obtain the state estimate in the original, physically meaningful coordinates is also a difficult problem [3]. This highlights the need for novel approaches that can overcome these limitations.
@@
In this paper, we present a novel learning approach for designing KKL observers for autonomous nonlinear systems. At the core of KKL observer design is an injective transformation that lifts the nonlinear system to a higher-dimensional space. The transformed system exhibits two key properties: its dynamics is linear up to output injection and is bounded-input-bounded-state stable. The KKL observer operates by replicating the transformed system in the higher-dimensional space. To reconstruct the state estimate in the original coordinates, the left-inverse of the transformation map is applied to the observer’s state. The transformation map’s injectivity property guarantees the accuracy of the state estimate.
@@
Given a KKL observer in an appropriate high-dimensional space, the transformation map can be obtained by solving a specific partial differential equation (PDE). However, this PDE is computationally challenging to solve in practice, and finding the left-inverse of the transformation map in real time presents additional difficulties. To address these challenges, we develop a framework for learning these maps using synthetic data generated from system dynamics and sensor measurements. We employ physics-informed neural networks to directly integrate the PDE constraint into the learning process for the transformation map and its inverse.
@@
This paper makes several key contributions to the learning-based design and implementation of KKL observers. In addition to proposing a physics-informed learning method for designing KKL observers, we establish theoretical robustness guarantees of the learned KKL observer against approximation errors and system uncertainties. To optimize observer performance, we develop a systematic method for parameter selection that enhances robustness while preserving the learning characteristics of the observer. Through extensive numerical simulations, we validate the effectiveness of our learning-based approach to KKL observer design. Additionally, we extend the practical utility of our method by demonstrating its application to sensor fault detection and isolation.
@@
The rest of the paper is organized as follows. Section 2 provides background on KKL observers and situates our work within the existing literature. Section 3 reviews the theoretical foundations of KKL observers. In Section 4, we present our novel learning method for KKL observer design, followed by Section 5 which develops theoretical robustness guarantees. Section 6 introduces our methodology for optimizing KKL observer parameters to find a trade-off between robustness and learnability. The effectiveness of our approach is demonstrated through numerical examples in Section 7, and with an application to fault detection and isolation presented in Section 8. Finally, Section 9 concludes the paper with a discussion of implications and future research directions.
@@ 2 Background and Related Literature
KKL observers generalize the theory of Luenberger observers [34, 35, 36] to nonlinear systems. Although the idea was initially proposed in [48] and [49], KKL observers were subsequently rediscovered by Kazantis and Kravaris [29], who provided local guarantees around an equilibrium point of the estimation error dynamics via Lyapunov’s Auxiliary Theorem. Later, [32] relaxed the restrictive assumptions of [29] to some extent; however, the analysis remained local until [31] proposed the first global result under the assumption of so-called finite complexity, which also turned out to be quite restrictive for general nonlinear systems. Subsequently, Andrieu and Praly [4] provided a comprehensive treatment of the KKL observer problem in 2006. Their key contribution lies in relating the existence of an injective transformation map – a requirement for KKL observers – to an observability-like property of the system known as backward distinguishability. In addition, Andrieu [2] proved that KKL observers converge exponentially and are tunable if the system is also differentially observable. The existence conditions are further refined in [12] and KKL observers with contracting dynamics are developed in [42]. When the observability conditions are not satisfied, [10] proposes a set-valued KKL observer that estimates a set of possible state trajectories that could have produced the observed output trajectories. Finally, extensions of KKL observers to non-autonomous and controlled nonlinear systems are presented in [21, 6, 8]; however, such systems are out of the scope of our present paper.
@@
The main challenge in synthesizing KKL observers for autonomous nonlinear systems is not only finding the transformation map that puts the system into a normal form but also finding an inverse map. Both problems turn out to be challenging in practice. To this end, [46, 44, 13, 41, 38, 45] have proposed several methods to approximate the transformation map and its inverse via deep neural networks.
@@
By fixing the dynamics of the KKL observer, [46] proposes to generate synthetic data trajectories by numerically integrating both the system model and the KKL observer, where both are initialized at multiple points in their corresponding state spaces. Then, using a supervised learning approach, a neural network is trained on the synthetic data to approximate the transformation map and its left-inverse. For discrete-time nonlinear systems, [44] proposes an unsupervised learning approach that enables proper exploration of the state space during training, whereas [45] extends a similar approach by allowing for switching observers. However, noisy measurements can lead to erroneous switching decisions, causing the observer to select the wrong observer mode or introduce instability due to chattering. Similarly, [13] proposes another unsupervised learning approach for tuning a KKL observer while adding the PDE associated with the transformation map as a design constraint. The unsupervised approach exhibits poor generalization capabilities. Specifically, when the real system’s initial state deviates significantly from the conditions observed during training, the observer’s performance deteriorates considerably. When the system dynamics are partially or fully unknown, [38] proposes a neural ODE-based approach to design KKL observers, which allows analyzing the trade-off between convergence speed and robustness, leading to improved observer training for robust performance.
@@
Our prior work [41] leveraged physics-informed learning to design KKL observers with improved accuracy, generalization, and training efficiency. In contrast to other learning-based approaches, we show that our method avoids overfitting and achieves better generalization performance across the entire state space. However, [41] employed a joint encoder-decoder neural network architecture to simultaneously learn the transformation map and its inverse, resulting in conflicting gradients between different components of the objective function. This caused the optimization to sometimes get stuck in a bad local minimum, which degraded the observer’s accuracy under certain cases [45]. In our current work, one of our contributions is to introduce a sequential learning approach for the KKL observer. We first learn the transformation map and subsequently utilize it to learn its inverse. This sequential approach significantly improves the observer’s accuracy.
@@ 3.1 Notation
For $v\in\mathbb{R}^{n}$, $\|v\|$ denotes its Euclidean norm and $\|v\|_{\infty}$ its max norm. For a matrix $M\in\mathbb{R}^{m\times n}$, $\|M\|$ denotes its spectral norm and $\text{cond}(M)$ its condition number. For a square matrix $M\in\mathbb{R}^{n\times n}$, $\text{eig}(M)\subset\mathbb{C}$ denotes the set of its eigenvalues and $\lambda_{\min}(M)\in\min_{\lambda\in\text{eig}(M)}|\lambda|$ denotes an eigenvalue of $M$ closest to the origin. We denote the set $[p]\coloneqq\{1,\dots,p\}$ for some $p\in\mathbb{Z}_{>0}$. Given a signal $s:\mathbb{R}_{\geq 0}\to\mathbb{R}^{n}$, we denote its restriction to $[0,t]$ by $s_{[0,t]}$, where $t\in\mathbb{R}_{\geq 0}$. Moreover, its essential supremum (or $L^{\infty}$) norm is defined as $\|s_{[0,t]}\|_{L^{\infty}}\coloneqq\inf\{c\in\mathbb{R}_{\geq 0}:\|s(\tau)\|_{\infty}\leq c~{}\text{for almost all}~{}\tau\in[0,t]\}$.
@@ 3.3 KKL Observer
where $A\in\mathbb{R}^{n_{z}\times n_{z}}$ and $B\in\mathbb{R}^{n_{z}\times n_{y}}$ are chosen such that $A$ is Hurwitz and the pair $(A,B)$ is controllable222$(A,B)$ is controllable iff $\text{rank}\left[\begin{array}[]{cccc}B&AB&\dots&A^{n_{z}-1}B\end{array}\right]=n_{z}$.. Although the transformed system (5) in the new coordinates is bounded-input bounded-state stable, the original nonlinear system (1) need not be Lyapunov stable. Moreover, (5) is not a linear system, but its linearity is only up to output injection, i.e., it is linear when the injected output $y=h(x)=h(\mathcal{T}^{*}(z))$ is ignored, where $\mathcal{T}^{*}$ is the left-inverse of $\mathcal{T}$.
Since
$\dot{z}(t)=\frac{\partial\mathcal{T}}{\partial x}(x(t))\dot{x}(t)$
it follows from (5) that, for a given $A\in\mathbb{R}^{n_{z}\times n_{z}}$ and $B\in\mathbb{R}^{n_{z}\times n_{y}}$, $\mathcal{T}$ must satisfy the following PDE:
@@
To obtain a state estimate $\hat{x}(t)$ in the original coordinates $\mathbb{R}^{n_{x}}$, the map $\mathcal{T}$ must be injective, which implies the existence of its left-inverse $\mathcal{T}^{*}$, i.e., $\mathcal{T}^{*}(\mathcal{T}(x))=x$. When the inverse exists, the KKL observer is obtained as
@@
where $\hat{z}(t)$ denotes the estimate of $z(t)$. Notice that KKL observer (7) is a special case of (2), where $\Phi(\hat{z},y)$ is affine in $\hat{z}$ and $\Psi(\hat{z},y)\coloneqq\mathcal{T}^{*}(\hat{z})$.
@@ 3.4 Existence Results
Let $x(t;x_{0})$ denote the solution trajectory of (1a) initialized at $x_{0}\in\mathcal{X}$.
@@
This assumption restricts our attention to nonlinear systems whose state $x(t)$ remains bounded in forward time.
@@
$\|z-\hat{z}\|\leq\varrho(\|\mathcal{T}^{*}(z)-\mathcal{T}^{*}(\hat{z})\|).$ (9)
@@
For the existence of a KKL observer (7) satisfying the estimation error requirement (4), it is sufficient that (1) is forward complete and the map $\mathcal{T}$ satisfying the PDE (6) is uniformly injective; see [4, Theorem 1]. $\diamond$
@@
Theorem 7 (Brivadis et al. [12]).
Let Assumptions 2 and 6 hold. Then, for almost any $(A,B)\in(\mathbb{R}^{n_{z}\times n_{z}},\mathbb{R}^{n_{z}\times n_{y}})$ such that
@@
Theorem 7 shows that there exists an injective map $\mathcal{T}$ and its left-inverse $\mathcal{T}^{*}$ such that the estimate $\hat{x}(t)$ obtained from the KKL observer (7) asymptotically converges to the true state $x(t)$. Therefore, relying on Assumptions 2 and 6, and fixing $n_{z}=n_{y}(2n_{x}+1)$, $A\in\mathbb{R}^{n_{z}\times n_{z}}$ a Hurwitz matrix, and $B\in\mathbb{R}^{n_{z}\times n_{y}}$ such that $(A,B)$ is controllable, we can learn the transformation map $\mathcal{T}$ and its left-inverse $\mathcal{T}^{*}$.

### v2

@@ Abstract
This paper proposes a novel learning approach for designing Kazantzis-Kravaris/Luenberger (KKL) observers for autonomous nonlinear systems. The design of a KKL observer involves finding an injective map that transforms the system state into a higher-dimensional observer state, whose dynamics is linear and stable. The observer’s state is then mapped back to the original system coordinates via the inverse map to obtain the state estimate. However, finding this transformation and its inverse is quite challenging. We propose learning the forward mapping using a physics-informed neural network, and then learning its inverse mapping with a conventional feedforward neural network. Theoretical guarantees for the robustness of state estimation against approximation error and system uncertainties are provided, including non-asymptotic learning guarantees that link approximation quality to finite sample sizes. The effectiveness of the proposed approach is demonstrated through numerical simulations on benchmark examples, showing superior generalization capability outside the training domain compared to state-of-the-art methods.
@@
thanks: This work is supported by the European Union’s Horizon Research and Innovation Programme under Marie Skłodowska-Curie grant agreement No. 101062523. It has also received funding from the Swedish Research Council and the Knut and Alice Wallenberg Foundation, Sweden. Corresponding author: M. U. B. Niazi.
@@ 1 Introduction
Observers play a crucial role in applications such as output feedback control [49], fault diagnosis [18], and digital twins [19]. Accurate state estimation is critical for effective control, monitoring, and decision-making in various engineering domains. Full-state measurement of real-world systems is often impractical, or even impossible, due to limitations in sensing resources and capabilities. This necessitates the use of observers that employ available sensor measurements and the system’s model to estimate the state.
A variety of observer designs exist, including Luenberger-like observers, extended Kalman filters, and high-gain observers, each with its own strengths and weaknesses depending on the specific application and system characteristics [24, 28, 9, 13]. This work focuses on Kazantzis-Kravaris/Luenberger (KKL) observers due to their applicability to a very general class of nonlinear systems and the well-developed theoretical framework for their design and analysis [27, 4, 2, 11, 14].
@@
Despite the advancements, synthesizing KKL observers remains a challenging task. The existing techniques assume the knowledge of an injective map that transforms a general nonlinear system into a special form required by the KKL observer. However, analytically finding this map is inherently difficult. Moreover, when this transformation map is known [10], finding its inverse to obtain the state estimate in the original, physically meaningful coordinates is also a difficult problem [3]. This highlights the need for novel approaches that can overcome these limitations.
@@
In this paper, we present a novel learning approach for designing KKL observers for autonomous nonlinear systems. At the core of KKL observer design is an injective coordinate transformation that lifts the nonlinear system to a higher-dimensional space. The transformed system exhibits two key properties: its dynamics is bounded-input-bounded-state stable and linear up to output injection. The KKL observer operates by replicating the transformed system in the higher-dimensional space. To reconstruct the state estimate in the original coordinates, the left inverse of the transformation map is applied to the observer’s state. The transformation map’s injectivity property guarantees the accuracy of the state estimate.
@@
Given a KKL observer in an appropriate high-dimensional space, the transformation map is a solution to a specific partial differential equation (PDE), [27, 4]. However, this PDE is computationally challenging to solve in practice, and finding the left inverse of the transformation map in real time is further challenging. To address these challenges, we develop a framework for learning these maps using synthetic data generated from system dynamics and sensor measurements. We employ physics-informed neural networks (PINNs) to directly integrate the PDE constraint into the learning process for the transformation map and its inverse.
@@
This paper makes several key contributions to the learning-based design and implementation of KKL observers. We propose a physics-informed learning method for designing KKL observers and establish theoretical learning guarantees for this approach. Additionally, we provide robustness guarantees for the learned KKL observer against approximation errors and system uncertainties. Through extensive numerical simulations, we validate the effectiveness of our learning-based approach to KKL observer design and provide comparisons with state-of-the-art approaches.
@@
The rest of the paper is organized as follows. Section 2 provides background and related literature on KKL observers and situates our work within the existing literature. Section 3 reviews the theoretical foundations of KKL observers. Section 4 reviews relevant notions in machine learning theory. In Section 5, we present our learning method for KKL observer design, followed by Sections 6 and 7, which develop theoretical learning and robustness guarantees, respectively. The effectiveness of our approach is demonstrated through numerical examples and comparisons in Section 8. Finally, Section 9 concludes the paper with a discussion of implications and future research directions.
@@ 2 Background and Related Literature
KKL observers generalize the theory of Luenberger observers [32, 33, 34] to nonlinear systems. Although the idea was initially proposed in [45] and [46], KKL observers were subsequently rediscovered by Kazantis and Kravaris [27], who provided local guarantees around an equilibrium point of the estimation error dynamics via Lyapunov’s Auxiliary Theorem. Later, [30] relaxed the restrictive assumptions of [27] to some extent; however, the analysis remained local until [29] proposed the first global result under the assumption of so-called finite complexity, which also proved quite restrictive for general systems. Subsequently, Andrieu and Praly [4] provided a comprehensive treatment of the KKL observer problem. Their key contribution lies in relating the existence of an injective KKL transformation map to an observability-like property of the system known as backward distinguishability. In addition, [2] proved that KKL observers converge exponentially and are tunable if the system is also differentially observable. The existence conditions are further refined in [14], and KKL observers with contracting dynamics are developed in [40]. When the observability conditions are not satisfied, [12] proposes a set-valued KKL observer that estimates a set of possible state trajectories that could have produced the observed output trajectories. Finally, extensions of KKL observers to non-autonomous and controlled nonlinear systems are presented in [22, 8, 10]; however, such systems are beyond the scope of the present paper.
@@
The main challenge in synthesizing KKL observers for autonomous nonlinear systems is not only finding the transformation map that puts the system into a normal form, but also finding its inverse map. Both problems turn out to be challenging in practice. To this end, [44, 41, 15, 39, 36, 50, 42] have proposed several methods to approximate the transformation map and its inverse via deep neural networks.
@@
By fixing the dynamics of the KKL observer, [44] proposes generating synthetic data trajectories by numerically integrating the system model and the KKL observer, both initialized at multiple points in their respective state spaces. Then, using a supervised learning approach, a neural network is trained on the synthetic data to approximate the transformation map and its left inverse. For discrete-time nonlinear systems, [41] proposes an unsupervised learning approach that enables proper exploration of the state space during training, whereas [42] extends this approach by allowing switching observers. However, noisy measurements can lead to erroneous switching decisions, causing the observer to select the wrong observer mode or introduce instability due to chattering. Similarly, [15] proposes another unsupervised learning approach for tuning a KKL observer while adding the PDE associated with the transformation map as a design constraint. The unsupervised approach exhibits poor generalization capabilities. Specifically, when the real system’s initial state deviates significantly from the conditions observed during training, the observer’s performance deteriorates considerably. When the system dynamics are partially or fully unknown, [36] proposes a neural ODE-based approach to design KKL observers, enabling analysis of the trade-off between convergence speed and robustness and leading to improved observer training for robust performance.
@@
Our prior work [39] leveraged physics-informed learning to design KKL observers with improved generalization and training efficiency, and [16] used this approach for detecting and isolating sensor faults. In contrast to other learning-based approaches, we demonstrated that our method avoids overfitting and achieves better generalization performance across the entire state space. However, [39] employed a joint encoder-decoder neural network architecture to simultaneously learn the transformation map and its inverse, leading to conflicting gradients across components of the objective function. This causes the optimization to sometimes get stuck in a bad local minimum, which degrades the observer’s accuracy in some instances [42]. In our current work, we introduce a sequential learning approach for the KKL observer. We first learn the transformation map and subsequently utilize it to learn its inverse. This approach improves the observer’s accuracy by avoiding conflicting gradients during training. Furthermore, we provide theoretical non-asymptotic generalization bounds for the learned maps and prove that the resulting state estimation error is input-to-state stable with respect to approximation errors, model uncertainties, and measurement noise.
@@ 3.1 Notation
For $\in\mathbb{R}^{n}, v\|v\|$ denotes its Euclidean norm and $\|v\|_{\infty}$ its max norm. For a matrix $\in\mathbb{R}^{m\times n}, M\|M\|$ denotes its spectral norm and $\operatorname*{cond}(M)$ its condition number. For a square matrix $M\in\mathbb{R}^{n\times n}$, $\operatorname*{eig}(M)\subset\mathbb{C}$ denotes the set of its eigenvalues and $\lambda_{\min}(M)\in\min_{\lambda\in\operatorname*{eig}(M)}|\text{Re}(\lambda)|$ denotes an eigenvalue of $M$ closest to the imaginary axis. We denote the set $p]\coloneqq\{1,\dots,p\}$ for some $p\in\mathbb{Z}_{>0}$. Given a signal $s:\mathbb{R}_{\geq 0}\to\mathbb{R}^{n}$, we denote its restriction to $[0,t]$ by $s_{[0,t]}$, where $t\in\mathbb{R}_{\geq 0}$. Moreover, its essential supremum (or $L_{\infty}$) norm is defined as $\|s_{[0,t]}\|_{L_{\infty}}\coloneqq\inf\{c\in\mathbb{R}_{\geq 0}:\|s(\tau)\|_{\infty}\leq c~\text{for almost all}~\tau\in[0,t]\}$.
@@ 3.3 KKL Observer
where $A\in\mathbb{R}^{n_{z}\times n_{z}}$ and $B\in\mathbb{R}^{n_{z}\times n_{y}}$ are chosen such that $A$ is Hurwitz and the pair $(A,B)$ is controllable222(A,B)(A,B) is controllable iff $\operatorname*{rank}\left[\begin{array}[]{cccc}B&AB&\dots&A^{n_{z}-1}B\end{array}\right]=n_{z}$.. Although the transformed system (5) in the new coordinates is stable, the original nonlinear system (1) need not be (Lyapunov) stable. Moreover, (5) is not a linear system, but its linearity is only up to output injection, i.e., it is linear when the injected output $y=h(x)=h(\mathcal{T}^{*}(z))$ is ignored, where $\mathcal{T}^{*}$ is the left-inverse of $\mathcal{T}$.
Since $\dot{z}(t)=\frac{\partial\mathcal{T}}{\partial x}(x(t))\dot{x}(t)$, it follows from (5) that, for a given AA and $B$, $\mathcal{T}$ must satisfy the following PDE:
@@
To obtain a state estimate $\hat{x}(t)$ in the original coordinates $\mathbb{R}^{n_{x}}$, the map $\mathcal{T}$ must be injective, implying the existence of its left inverse $\mathcal{T}^{*}$, i.e., $\mathcal{T}^{*}(\mathcal{T}(x))=x$. When this inverse exists, the KKL observer is obtained as
@@
[…]
@@ 3.4 Existence Results
Let the solution of (1a) initialized at $x_{0}\in\mathcal{X}$ be $x(t;x_{0})$.
@@
This assumption restricts our attention to nonlinear systems whose state $x(t)$ remains in a bounded domain.
@@
$).\|z-\hat{z}\|\leq\varrho(\|\mathcal{T}^{*}(z)-\mathcal{T}^{*}(\hat{z}$$\|)$. (9)
@@
For the existence of a KKL observer (7) satisfying the requirement (4), it is sufficient that (1) is forward complete and the map $\mathcal{T}$ satisfying the PDE (6) is uniformly injective; see [4, Theorem 1]. $\diamond$
@@
Theorem 7.
(Brivadis et al. [14]). Let Assumptions 2 and 6 hold. Then, for almost any $(A,B)\in\mathbb{R}^{n_{z}\times n_{z}}\times\mathbb{R}^{n_{z}\times n_{y}}$ such that
@@
This theorem guarantees the existence of a uniformly injective map $\mathcal{T}$ and its left-inverse $\mathcal{T}^{*}$ such that the estimate $\hat{x}(t)$ obtained from the KKL observer (7) asymptotically converges to the true state $x(t)$. Therefore, relying on Assumptions 2 and 6, and fixing $n_{z}=n_{y}(2n_{x}+1)$, $A\in\mathbb{R}^{n_{z}\times n_{z}}$ a Hurwitz matrix, and $B\in\mathbb{R}^{n_{z}\times n_{y}}$ such that $(A,B)$ is controllable, we can learn the transformation map $\mathcal{T}$ and its left-inverse $\mathcal{T}^{*}$.
4 Preliminaries on Learning Theory
We revisit the standard framework of statistical learning theory, which is relevant to this paper.
4.1 Learning Problem
Consider a supervised learning problem where the goal is to learn a mapping $\hat{\mathcal{T}}:\mathcal{X}\to\mathcal{Z}$ from an input space $\mathcal{X}$ to an output space $\mathcal{Z}$. We assume an unknown, fixed probability distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Z}$. The quality of a prediction $\hat{\mathcal{T}}(x)$ for a true output $z$ is measured by a loss function $L(\hat{\mathcal{T}}(x),z)$.
The goal is to find a function $\hat{\mathcal{T}}$ that minimizes the true risk
$R(\hat{\mathcal{T}})=\mathbb{E}_{(x,z)\sim\mathcal{D}}[L(\hat{\mathcal{T}}(x),z)]$
[…]
$\hat{R}_{S}(\hat{\mathcal{T}})=\frac{1}{N}\sum_{i=1}^{N}L(\hat{\mathcal{T}}(x_{i}),z_{i})$
on this sample $S$.
4.2 Hypothesis Class and Error Decomposition
A learning algorithm searches for a solution within a predefined hypothesis class $\mathcal{H}$, which is a set of functions $\hat{\mathcal{T}}:\mathcal{X}\to\mathcal{Z}$. For example, $\mathcal{H}$ can be the set of all functions representable by a specific neural network architecture. The empirical risk minimizer is the hypothesis $\hat{\mathcal{T}}_{S}\in\mathcal{H}$ that minimizes the empirical risk, i.e.,
[…]
The excess risk $R(\hat{\mathcal{T}}_{S})-R^{*}$ of the learned hypothesis, where $R^{*}$ is the Bayes risk, can be decomposed into two fundamental components: approximation error
$\inf_{\hat{\mathcal{T}}\in\mathcal{H}}R(\hat{\mathcal{T}})-R^{*}$
and estimation error
[…]
The approximation error is deterministic and measures how well the chosen class $\mathcal{H}$ can, in principle, approximate the true optimal (Bayes) predictor. It is independent of the data size $N$. On the other hand, the estimation error is random (it depends on the sample $S$) and measures the penalty for having only a finite sample of $N$ data points instead of the true distribution $\mathcal{D}$. A central goal of learning theory is to bound the estimation error, which requires a measure of the richness (or complexity) of the hypothesis class $\mathcal{H}$, as defined in the following subsection.
4.3 Pseudo-Dimension and Generalization Bounds
For real-valued functions, such as those in regression or in our KKL observer learning problem, the VC-dimension [51] (used for binary classification) is extended to the pseudo-dimension [43, 25]. This concept is defined not on the hypothesis class $\mathcal{H}$ itself, but on the associated loss class $\mathcal{G}=\{(x,z)\mapsto L(\hat{\mathcal{T}}(x),z):\hat{\mathcal{T}}\in\mathcal{H}\}$.
Definition 8.
(Shattering and Pseudo-dimension [38]). A set of $d$ points $\{(x_{i},z_{i})\}_{i=1}^{d}$ is shattered by the loss class $\mathcal{G}$ if there exist thresholds $\tau_{1},\dots,\tau_{d}\in\mathbb{R}$ such that for all $2^{d}$ possible binary labelings $b\in\{-1,+1\}^{d}$, there exists $\mathcal{T}\in\mathcal{H}$ whose corresponding loss $g\in\mathcal{G}$ satisfies
$\forall i\in\{1,\dots,d\},\quad\mathrm{sign}(g(x_{i},z_{i})-\tau_{i})=b_{i}.$
The pseudo-dimension of $\mathcal{G}$, denoted $\mathrm{Pdim}(\mathcal{G})$, is the size $d$ of the largest set that can be shattered by $\mathcal{G}$.
The pseudo-dimension provides a combinatorial measure of the complexity of $\mathcal{H}$ via its loss class $\mathcal{G}$. A finite pseudo-dimension $d$ ensures that the estimation error will decrease as the sample size $N$ increases. This is formalized in a generalization bound below.
Lemma 9.
(Generalization Bound for Bounded Regression [38, Theorem 11.8]). Let $\mathcal{G}$ be a loss class with pseudo-dimension $d=\mathrm{Pdim}(\mathcal{G})<\infty$. Assume the loss $L$ is bounded by a constant $M$. Then, for any $\delta>0$, with probability at least $1-\delta$ over the draw of an i.i.d. sample $S$ of size $N$, it holds that for every $\hat{\mathcal{T}}\in\mathcal{H}$,
$R(\hat{\mathcal{T}})\leq\hat{R}_{S}(\hat{\mathcal{T}})+M\sqrt{\frac{2d\ln(eN/d)}{N}}+M\sqrt{\frac{\ln(1/\delta)}{2N}}.$
This lemma provides an explicit, non-asymptotic bound on the estimation error. The complexity term
$\mathcal{C}(\mathcal{H},N,\delta)\triangleq M\sqrt{\frac{2d\ln(eN/d)}{N}}+M\sqrt{\frac{\ln(1/\delta)}{2N}}$
depends on the hypothesis class $\mathcal{H}$ via $d$ and the sample size $N$. For the learning procedure of KKL observers developed in this paper, we derive such explicit non-asymptotic error bounds.